A woman driving at the speed limit in a 25 mi/hour zone enters a zone where the speed limit is 45 mi/hour. She accelerates at a constant rate and reaches the new speed limit in 6.00 s. What distance does the car travel during that acceleration? (Be careful with your units) 1.
Answer #1

V = 45mi/hr = 20.12m/s

U = 25mi/hr = 11.17m/s

From V - U = at

=) a = (20.12 - 11.17)/6 = 1.404 m/s^2

Now distance S = (V^2 - U^2 )/2a = (20.12^2 - 11.17^2 )/2×1.404

=) S = 97.6m